Suppose you recieve an inheritance of 5000 and invest it at 5% interest, compounded continuously. The equation C(x) = 5000e^0.05x can be used to model this investment, where x is the number of years your money is invested and C(x) represents the amount of money you have
A. How much Will you have in 5 years
B. How long Will it take for your money to double

Answer by rfer(16322)

FV=5000(2.7183)^(0.05*5)
FV=5000(2.1783)^0.25
FV=5000(1.2148689)
FV=6074.35